OpenStudy (anonymous):

find the slope of the curve y = x^2 (x+2)^2 at point (1,2) a. 81 b. 48 c. 64 d. 54

OpenStudy (anonymous):

take the derivative, replace x by 1

OpenStudy (anonymous):

my answer is 64 i got wrong... i just wanted to know the answer...

OpenStudy (anonymous):

\[x^2(x+2)^2=x^2(x^2+4x+4)=x^4+4x^3+4x\] \[y'=4x^3+12x^2+4\] i get 20

OpenStudy (anonymous):

the answer is in the letter i just forgot..and i need it

OpenStudy (anonymous):

oh i made a mistak

OpenStudy (anonymous):

\[x^2(x+2)^2=x^2(x^2+4x+4)=x^4+4x^3+4x^2\] \[y'=4x^3+12x+8\] now i get 24

OpenStudy (anonymous):

i get 24 omg

OpenStudy (jlastino):

wth

OpenStudy (amistre64):

there is no point 1,2 on this curve

OpenStudy (anonymous):

OMG MY CLASSMATE KING

OpenStudy (anonymous):

KINGkos is review his quiz and exam too hahahahha

OpenStudy (jlastino):

so the truth is revealed haha

OpenStudy (anonymous):

on tuesday is final exam thats y i memorized all the answer in quiz and homework.this question is my X in the quiz

OpenStudy (amistre64):

your question makes no sense ... chk it for typos

OpenStudy (anonymous):

http://openstudy.com/study#/updates/4f8a50eae4b09e61bffc0287 correct question